# A Morphology-Based Investigation of Positional Encodings

### v1

@@ Abstract
How does the importance of positional encoding in pre-trained language models (PLMs) vary across languages with different morphological complexity? In this paper, we offer the first study addressing this question, encompassing 23 morphologically diverse languages and 5 different downstream tasks. We choose two categories of tasks: syntactic tasks (part-of-speech tagging, named entity recognition, dependency parsing) and semantic tasks (natural language inference, paraphrasing). We consider language-specific BERT models trained on monolingual corpus for our investigation. The main experiment consists of nullifying the effect of positional encoding during fine-tuning and investigating its impact across various tasks and languages. We perform additional experiments on a subset of tasks and languages, where we deliberately distort the original word order of a sentence. Our findings demonstrate that the significance of positional encoding diminishes as the morphological complexity of a language increases. Across all experiments, we observe clustering of languages according to their morphological typology - with analytic languages at one end and synthetic languages at the opposite end.
Keywords: Pre-trained language models (PLMs), Positional Encoding, Morphosyntax, Linguistic Diversity
[…]
@@ A Morphology-Based Investigation of Positional Encodings
Poulami ${}^{\clubsuit}$, Shikhar ${}^{\spadesuit}$, Raj ${}^{\diamondsuit}$, Pushpak ${}^{\clubsuit}$
${}^{\clubsuit}$omputer Science and Engineering, IIT Bombay, India
${}^{\spadesuit}$oogle Research, India ${}^{\diamondsuit}$, Japan
{poulami, pb}@cse.iitb.ac.in
Abstract content
1.   Introduction
Transformers Vaswani et al. (2017) have ushered in a new era in the field of natural language processing, facilitating the development of highly versatile language models (LMs) that have achieved ground-breaking results across a wide spectrum of language processing tasks such as natural language inference Liu et al. (2019a), text classification Raffel et al. (2020a), named entity recognition Liu et al. (2019a), and part-of-speech tagging Martin et al. (2020). This has led to the emergence of various pre-trained language models Lan et al. (2019); Xue et al. (2020); Brown et al. (2020b). The training of these models on vast quantities of raw text equips them with the ability to accurately capture the inherent linguistic features of the language Tenney et al. (2019). The general paradigm is to further fine-tune a pre-trained model on task-specific or domain-specific data for any downstream task.
However, existing models often overlook the diverse fine-grained and broader domains of linguistics. This is particularly evident since there has been very few instances where morphology of a language has been taken into consideration while designing a language model. For instance, Nzeyimana and Rubungo (2022) proposes a morphology-aware language model build on a two-tier transformer architecture for low-resource morphologically rich Kinyarwanda language. Park et al. (2021) study the impact of morphological complexity on LSTM based language models. As language models are being developed for more and more languages which significantly differ in morphological typology, it is imperative to incorporate linguistic knowledge and construct language models that are sensitive to these linguistic nuances. This serves as the motivation behind our work in this paper.
Figure 1: Methodology Overview
Figure 2: The figure illustrates the effect of word order on semantics for two languages: English (left) and Sanskrit (right). English is a morphologically poor language with SVO word order whereas Sanskrit is a morphologically rich language with no dominant word order(NODOM). Distorting the word order completely alters the meaning for English. However, for Sanskrit the meaning remains intact. More details in Section 4.1.
The vast diversity of morphological systems exhibited in languages worldwide constitutes an intriguing dimension of human language. The central focus of our paper revolves around investigating the relationship between morphology and word order by leveraging the concept of positional encoding. Our contributions are:
@@ 1.   Introduction
  • We perform the first study about the varying importance of positional encoding across languages with different morphological complexity.
@@
  • We are the first to show that the impact of positional encodings diminishes as the morphological complexity of a language increases.
@@
  • We conduct exhaustive experiments covering 23 different languages across 8 language families and 5 diverse natural language processing tasks.
  • Across all results, we observe clustering of languages according to their morphological typology with analytic languages at one end and synthetic languages at the opposite end.
2.   Related work
In this section, we provide a comprehensive overview of pre-trained language models, positional encodings and how they relate to linguistics.
Pre-trained language models: Pre-trained language models such as BERT(Devlin et al., 2018), RoBERTa(Liu et al., 2019b), GPT-2(Radford et al., 2019), T5(Raffel et al., 2020b), GPT-3(Brown et al., 2020a) have led to significant advancements in the field of natural language processing. However, the enormous computational cost incurred during their training is a major challenge in the development of pre-trained language models. Acquiring a deeper understanding of how various components of a pre-trained language model function in different languages can provide valuable insights regarding their necessity (or redundancy) across languages, helping in the efficient reduction of computational costs.
Positional Encodings: Positional encoding is a crucial component in transformer-based language models. It is used to explicitly add positional information and capture word order within a sequence. There exist several variants of positional embeddings such as absolute positional embeddings Vaswani et al. (2017); Radford et al. (2019), and relative positional embeddings Shaw et al. (2018); Raffel et al. (2020b); Devlin et al. (2018). They were originally introduced for English, a morphologically poor language. Thus, it is uncertain whether their effectiveness can be generalized across languages with varying morphological complexity.
Probing pre-trained language models for linguistic knowledge: As PLMs are being developed for more and more languages that significantly differ in linguistic complexity, it is imperative to incorporate linguistic knowledge and construct language models that are sensitive to different linguistic nuances. Recently, there has been a surge in the study of linguistic knowledge encoded by pre-trained language models, focusing different aspects of linguistics such as syntax (Goldberg, 2019; Jawahar et al., 2019) and semantics (Ethayarajh, 2019). There has also been several studies to understand how morphology is encoded in PLMs. Sadeniemi et al. (2008) investigate the morphological, morpho-syntactic and syntactic complexities of European Union languages by using a compression based metric motivated by Kolmogorov complexity. Gerz et al. (2018) investigate the connection between language modeling and linguistic typology across 50 different languages. However their work do not include transformer-based language models and encompass a limited set of tasks. Edmiston (2020) probe BERT models of five European languages for morphological information. Otmakhova et al. (2022) analyze how different layers of the Transformer encode morphology and syntax. But they focus their study only on three languages and devise a simple classification task such as identification of word order corruption to understand the correlation between word order and morphology in BERT.
Infusing linguistic knowledge in language models: Over the past few years, there has been significant efforts to infuse linguistic knowledge into pre-trained language models. Chaudhary et al. (2018) adapt word embeddings by leveraging morphological and phonological information. Nzeyimana and Niyongabo Rubungo (2022) employed feature-based pre-training to design a morphology-aware language model for a low-resource African language, Kinyarwanda. Chakrabarty et al. (2022) propose a encoder-decoder pre-trained model enriched with linguistic features such as lemma, part-of-speech and dependency labels. All these works show that incorporating linguistic knowledge improves the performance of pre-trained models in downstream tasks. However, they don’t shed light on how the relevance of different components of a transformer model can vary across languages.
3.   The Relationship between Morphology and Word Order
@@
Linguistic theories : Morphological case markings serve a similar function as word order(Sapir, 1921; Blake, 2001).Several theories suggest that the presence of morphological case is necessary for free word order in a language (Comrie, 1989; Haspelmath, 1999). Either morphological case or structural position facilitates the unambiguous determination of grammatical role of the constituents of a sentence. The existence of morphological case reduces the need for fixed structural position in determination of the grammatical function of a word or phrase, allowing for variable word order. However, if morphological case is absent, fixed placement of words (and phrases) is necessary, exhibiting a fixed or rigid word order. In this work, we align our empirical study in accordance with the above theories that hints at the existence of a correlation between morphology and word order. Specifically, morphologically rich languages which tend to exhibit higher word-order flexibility as compared to morphologically poor languages [^1] The theory regarding morphology and word order is a linguistically complex topic. Concurrent theories in the literature propose that there is no correlation between morphological complexity and word order. Müller (2002) demonstrates phenomena like scrambling and topic shift, where the change in word order does not necessarily require a high level of morphological complexity. However, this is beyond the scope of our study..
3.1.   Spectrum of Morphological Complexity
@@
4.   Methodology
4.1.   Morphology-based Investigation
In Transformer-based Vaswani et al. (2017) language models such as BERT Devlin et al. (2018), a given input sequence $S$ consisting of tokens $w_{1},w_{2},...w_{n}$ is fed to an embedding layer to generate the corresponding word embeddings $\bm{e}_{1},\bm{e}_{2},...,\bm{e}_{n}$. This is followed by multiple self-attention layers. However, the attention mechanism is completely invariant to the order of tokens which is crucial for capturing semantics in languages such as English. Figure 2 illustrates that how the same set of tokens in English can have completely different meaning given their different ordering. Hence, positional embeddings are added to the word embeddings to effectively capture the ordering of tokens. For each token $w_{i}$, a real-valued vector $\bm{p}_{i}\in\mathbb{R}^{d}$ is added to its token embedding $\bm{e}_{i}$. Thus, a position-aware token embedding $\bm{e}^{\prime}_{i}=\bm{e}_{i}+\bm{p}_{i}$ is fed to the self-attention stack. Vaswani et al. (2017) proposes the use of sine and cosine functions of different frequencies to encode positional information.
$PE_{(pos,2i)}=sin\left(\frac{pos}{10000^{\frac{2i}{d_{model}}}}\right)$
$PE_{(pos,2i+1)}=cos\left(\frac{pos}{10000^{\frac{2i}{d_{model}}}}\right)$
where $pos$ represents the position of a token in a sequence and $i$ denotes the $i^{th}$ dimension of the embedding vector. To study the impact of positional embeddings on Transformer-based LMs for languages with different morphological richness, we propose the following two strategies.
  • Nullifying Positional Encoding: In this strategy, we set positional embedding in the language model to null, i.e., $\bm{p}_{i}=\bm{0}$. Thus, making model invariant to the token word of the input sequence. We hypothesize that for morphological rich languages such as Sanskrit this would not have much impact on the downstream performance. For instance, in Figure 2, we demonstrate that for Sanskrit, the meaning of the sentence remains consistent even when the token order is shuffled. However, this does not hold for morphologically poor languages such as English.
  • Permuting Word Order: Here, we distort the word order by shuffling a sentence at the level of uni-gram, bi-gram and tri-gram. The primary objective is to introduce varying degrees of distortion and examine the resulting loss in information across morphologically diverse languages. The proposed permutation strategy can be applied to study the robustness of the language models to the token order of the input.
5.   Experimental Setup
To ensure the generalizability of our findings, we choose to perform a comprehensive study spanning different languages and tasks. In this section, we outline the principles for selecting diverse tasks and languages and provide details on the datasets and models used in evaluation.
5.1.   Tasks
The main objective of our research is to analyze the effect of positional encoding in languages with varying morphological complexity. As discussed in Section 2, positional encoding encodes word order and thereby the syntactic relationships inherent in a language. Since our work deals with the interplay of morphology and syntax in pre-trained language models, we consider two sets of tasks:
  • Syntactic tasks - Part-of-speech tagging, Named Entity Recognition, Dependency Parsing
  • Semantic tasks - Natural Language Inference(NLI), Paraphrasing
This choice of tasks enables us to understand the function of morphology and word order in encoding syntactic and semantic role information across various languages. Moreover, text classification tasks such as natural language inference and paraphrasing operate at the sentence level. On the other hand, part-of-speech tagging and named entity recognition are sequence labeling tasks that assess the model’s competence at a more granular word level. Additional factors considered in task selection include:
  • [itemsep=2pt,parsep=2pt,partopsep=2pt,leftmargin=*,topsep=2pt]
  • Multilinguality: We select tasks that span a diverse set of languages and language families to establish the broad applicability of our results.
  • Sufficient monolingual training data: We consider tasks with sufficient monolingual training data. As we remove positional encoding at the time of fine-tuning, the training data needs to be substantial for the impact of positional encoding removal to be profound.
5.2.   Languages
The choice of our languages is determined by the fulfillment of the following criteria: (1) availability of monolingual BERT-base model on HuggingFace Hub (Wolf et al., 2019), (2) availability of sufficient monolingual training data across different tasks and (3) typological diversity. We aim to cover as many languages and language families as possible. Overall, we cover 23 languages distributed across 8 language families and one language isolate. In Table 1, we present an outline of these languages
@@ 5.2.   Languages
Basque (eu) Basque (Language isolate) orai-nlp/ElhBERTeu
@@
Hindi (hi) Indo-European: Indo-Aryan l3cube-pune/hindi-bert-scratch
Hungarian (hu) Uralic SZTAKI-HLT/hubert-base-cc
@@
Malay (ms) Austronesian malay-huggingface/bert-base-bahasa-cased
Persian (fa) Indo-European: Iranian HooshvareLab/bert-base-parsbert-uncased
@@
5.3.   Datasets
In our experiments, we evaluate our models on a variety of tasks. We utilize GLUE Wang et al. (2018) benchmark which comprises of nine language understanding tasks encompassing a wide spectrum of domains. The benchmark includes sentence acceptability judgment on CoLA Warstadt et al. (2018), sentiment analysis on SST-2 Socher et al. (2013), paraphrasing/sentence similarity on MRPC Dolan and Brockett (2005), STS-B Cer et al. (2017), QQP Sharma et al. (2019), and natural language inference on MNLI Williams et al. (2018), QNLI Rajpurkar et al. (2016), and RTE Bentivogli et al. (2009)) datasets.
In order to broaden the scope of our analysis beyond the English language, we incorporate multilingual datasets that span various domains and encompass multiple languages. Our investigation includes tasks from the XTREME benchmark Hu et al. (2020), covering natural language inference (XNLI) Conneau et al. (2018), paraphrasing (PAWS-X) Yang et al. (2019), and structure prediction tasks such as part-of-speech (POS) tagging and named entity recognition (NER). We use the data from the Universal Dependencies v2.12 (Zeman et al., 2023) for the task of dependency parsing. The treebanks used for different languages are specified in Table 3. By incorporating these tasks, we aim to enhance the comprehensiveness and breadth of our research.
5.4.   Evalution Metrics
5.5.   Models
In this section, we describe the baseline and perturbed models used in our experiments.
5.6.   Model Selection
In our research, we explore monolingual pre-trained language models to prevent cross-lingual transfer from influencing our results. Given the availability of monolingual BERT models in various languages, we select BERT as the example PLM for our study. We consider BERT-base model for all languages to ensure that variations in model size and architecture do not influence the results.
5.6.1.   Baseline Models
We used fine-tuned BERT-base models as the baseline.
5.6.2.   Perturbed Models
We conduct experiments on two perturbed models, as outlined in Section 4.1.
  • Models without positional encoding: During the fine-tuning stage, we nullify positional encoding and measure the relative drop in performance with respect to the baseline model to assess how positional encoding impacts different languages. [^2] Training language models from scratch without positional encoding requires significant computational power and financial resources. Hence, due to limited computational resources, we focused on fine-tuning rather than pre-training. We discuss this further in section Limitations.
  • Models with perturbed input: Given a sentence, we compute a list of N-grams and randomly shuffle the entire list to generate the distorted sentence. We ensure that the shuffled sentence is different from the original sentence. We use non-perturbed training and development sets and only perturb the test inputs. We measure the relative drop in performance with respect to the baseline model to observe how word order corruption impacts different languages.
5.7.   Training and evaluation setup
For text classification tasks like natural language inference, paraphrasing and sequence labelling tasks like named-entity recognition and POS tagging, we adhered to the generic pipeline. We implemented a biaffine parser for dependency parsing, by applying a biaffine attention layer directly on the output of BERT as described in Glavaš and Vulić (2021). As outlined in Devlin et al. (2018), we have performed hyper-parameter tuning and fine-tuned BERT on GLUE tasks. For multilingual experiments, hyper-parameter tuning is performed on English validation data as suggested in the XTREME benchmark. However, since our goal is not to achieve the best absolute performance, we avoided conducting extensive hyperparameter tuning. For the reproducibility of our results, we will add the hyperparameter details in the appendix of the paper.

### v2

@@ Abstract
Contemporary deep learning models effectively handle languages with diverse morphology despite not being directly integrated into them. Morphology and word order are closely linked, with the latter incorporated into transformer-based models through positional encodings. This prompts a fundamental inquiry: Is there a correlation between the morphological complexity of a language and the utilization of positional encoding in pre-trained language models? In pursuit of an answer, we present the first study addressing this question, encompassing 22 languages and 5 downstream tasks. Our findings reveal that the importance of positional encoding diminishes with increasing morphological complexity in languages. Our study motivates the need for a deeper understanding of positional encoding, augmenting them to better reflect the different languages under consideration.
@@ A Morphology-Based Investigation of Positional Encodings
Poulami Ghosh, Shikhar Vashishth, Raj Dabre, Pushpak Bhattacharyya IIT Bombay, India, Google Research, India NICT, Japan {poulami, pb}@cse.iitb.ac.in
1 Introduction
Pre-trained language models (PLMs) Devlin et al. (2018); Liu et al. (2019a); Radford et al. (2019); Raffel et al. (2020a); Brown et al. (2020) built upon transformers Vaswani et al. (2017) have achieved ground-breaking results across a wide spectrum of language processing tasks such as natural language inference Liu et al. (2019b), text classification Raffel et al. (2020b), named entity recognition Liu et al. (2019b), and part-of-speech tagging Martin et al. (2020). However, only a few models take into account various linguistic aspects and theories in their design Nzeyimana and Rubungo (2022); Park et al. (2021). Morphology and word order of a language are closely related Sapir (1921); Comrie (1989); Blake (2001); the latter is incorporated into transformer-based models through positional encoding (PE) Dufter et al. (2022). As language models are being developed for more languages which significantly differ in morphological typology, it is could be beneficial to construct language models that are sensitive to these linguistic nuances. Moreover, the enormous computational cost incurred during their training is a major challenge in the development of PLMs. Acquiring a deeper understanding of how various components of a PLM function in different languages can provide valuable insights regarding their necessity across languages. This motivates us to investigate the relation between positional encoding and morphology, which is essential for wider usage of PLMs across different languages. Our contributions are:
@@ 1 Introduction
1. We perform the first study about the varying importance of positional encoding across languages with different morphological complexity.
@@
2. We show that the impact of PE diminishes as the morphological complexity of a language increases.
@@
3. We conduct exhaustive experiments covering 22 different languages across 9 language families and 5 diverse natural language processing tasks.
Figure 1: The figure illustrates the effect of word order on semantics for two languages: English (left) and Sanskrit (right). English is a morphologically poor language with SVO word order whereas Sanskrit is a morphologically rich language with no dominant word order (NODOM). Distorting the word order completely alters the meaning for English. However, for Sanskrit the meaning remains intact.
2 Related work
Positional Encoding (PE): Various methods have been proposed to incorporate position information in transformer models. Absolute positions in a sequence, represented by fixed Vaswani et al. (2017) or trainable encodings Gehring et al. (2017); Devlin et al. (2018); Radford et al. (2019); Lan et al. (2019), are typically added to input embeddings. Relative positions are encoded by directly adding position biases into the attention matrix Shaw et al. (2018); Yang et al. (2019b); Raffel et al. (2020a); Huang et al. (2020); He et al. (2020); Press et al. (2021). Su et al. (2021) introduce rotary positional embeddings, employing a rotation matrix to encode both absolute and relative position information. Our study focuses on BERT models, which use learnable absolute PEs Wang et al. (2020); Huang et al. (2020). We draw insights from linguistics theories and question the design choices for BERT-style models that were designed with English in mind.
Absence of Positional Information: Eliminating positional encoding results in a bag-of-words representation. Sinha et al. (2021) pre-train a RoBERTa model without positional embeddings and observe degraded performance on GLUE and PAWS tasks. Haviv et al. (2022); Le Scao et al. (2022) demonstrate that causal language models lacking explicit PE remain competitive with standard position-aware models. Additionally, Haviv et al. (2022) find that a pre-trained RoBERTa large model without PE exhibits higher perplexities than position-informed models. However, all these observations are limited to the English language.
Linguistic Information in BERT: Several works studied the linguistic knowledge encoded in PLMs such as BERT, focusing on different aspects of linguistics such as syntax (Goldberg, 2019; Jawahar et al., 2019), semantics (Ethayarajh, 2019) and morphology (Edmiston, 2020). Tenney et al. (2019); Puccetti et al. (2021) investigate the extent and organization of the linguistic information encoded in BERT. Gerz et al. (2018) investigate the connection between language modeling and linguistic typology across 50 different languages. However, they do not consider PLMs. Otmakhova et al. (2022) examine how various layers within a BERT model encode morphology and syntax.
3 Methodology
In our work, we first quantify morphological complexity, and then systematically study the effect of removal of positional encodings during fine-tuning. Please refer to Appendix A for details on linguistic theories governing our study.
3.1 Quantifying Morphological Complexity
Following Kettunen (2014); Jayanthi and Pratapa (2021); Çöltekin and Rama (2023), we employ type-token ratio (TTR) as an empirical proxy of morphological complexity of a language. We use the many-to-many multilingual Flores-200 benchmark Costa-jussà et al. (2022) to ensure information consistency across languages. As Chinese is an unsegmented language, we use character level ELMo model from the pywordseg library Chuang (2019) to split Chinese text into words. The remaining languages are space-delimited. Please refer to Appendix C for more details.
3.2 Morphology-based Investigation
To study the impact of positional embeddings, we set them to 0, effectively nullifying its effect during fine-tuning. We posit that for morphologically rich languages like Sanskrit, this would have minimal impact on downstream performance. For example, as depicted in Figure 1, the semantic meaning of a sentence in Sanskrit remains consistent even when the order of tokens is shuffled. However, this does not hold for morphologically poor languages.
4 Experimental Setup
To ensure the generalizability of our findings, we choose to perform a comprehensive study spanning different languages and tasks.
4.1 Tasks and Languages
As our work deals with the interplay of morphology and syntax in PLMs, we consider two sets of tasks: a. Syntactic tasks: Part-of-speech (POS) tagging, Named Entity Recognition (NER), Dependency Parsing b. Semantic tasks: Natural Language Inference(NLI), Paraphrasing
Factors considered in task and language selection include (1) availability of monolingual BERT-base model on HuggingFace Hub (Wolf et al., 2019), (2) availability of sufficient monolingual training data across different tasks, and (3) typological diversity. We aim to cover as many languages and language families as possible. Overall, we cover 22 languages distributed across 9 language families and one language isolate. We present an outline of the languages in Appendix B due to space constraints.
4.2 Datasets
Our study includes tasks from the XTREME benchmark Hu et al. (2020), covering natural language inference (XNLI) Conneau et al. (2018), paraphrasing (PAWS-X) Yang et al. (2019a), and structure prediction tasks such as POS tagging and NER. We use the data from the Universal Dependencies v2.12 (Zeman et al., 2023) for the task of dependency parsing. The treebanks used for different languages are listed in Table 4 in Appendix.
4.3 Model Selection
In our research, we use monolingual pre-trained language models to prevent cross-lingual transfer from influencing our results. Given the availability of monolingual BERT models in various languages, we select BERT as the example PLM for our study. We consider BERT-base model for all languages to ensure that variations in model size and architecture do not influence the results. We consider fine-tuned BERT-base models with PE and without PE as the baseline and perturbed models, respectively.
4.4 Evalution Metrics
The metric used for different tasks is outlined in Table 1. For a given task, let m and n denote the metric scores for the baseline and perturbed models, respectively. We use the relative decrease in performance, calculated as (m-n)/n, as a quantitative measure of the importance of PE on the language. A higher value indicates a greater utilization of PE in effectively modeling the language.
4.5 Training and evaluation setup
For text classification tasks, we follow the generic pipeline. For dependency parsing, we implemented a biaffine parser by applying a biaffine attention layer directly on the output of BERT as described in Glavaš and Vulić (2021). As suggested in the XTREME benchmark, we have performed hyper-parameter tuning on English validation data. However, since our goal is not to achieve the best absolute performance, we avoided conducting extensive hyperparameter tuning. More details are present in the Appendix F. Results are reported across 3 random trials of each experiment.
Figure 2: Effect of Positional Encoding on NER task.
Figure 3: Effect of Positional Encoding on POS task.
Figure 4: Effect of Positional Encoding on Dependency Parsing.
5 Results
In this section, we present the findings of our experiments on syntactic and semantic tasks. We also conducted preliminary experiments on the GLUE benchmark, as discussed in Appendix E.
5.1 Results on Syntactic Tasks
Figures 2, 3 demonstrate the effect of removing positional encoding in NER, POS tagging tasks. For dependency parsing, figures 4 and 7 depict the effects on UAS and LAS scores, respectively.
a. Analytic languages like Chinese and Vietnamese, characterized by minimal or no morphology, exhibit the most significant decrease in performance when PE is removed. Moderately analytic languages like English and French follow.
b. In synthetic languages such as Hungarian, Finnish, and Turkish, known for their rich morphological systems, the function of morphology in encoding grammatical roles surpasses that of word order, resulting in a considerably smaller decrease in performance when PE is eliminated.
Figure 5: Effect of Positional Encoding on XNLI.
Figure 6: Effect of Positional Encoding on PAWS-X.
In the XTREME benchmark, different methods were employed for data annotation for POS tagging and NER datasets. While the former was human annotated, the latter was created through automatic annotation using weak supervision. Despite these disparities, the findings of the POS tagging and NER experiments are similar.
5.1.1 Results on Semantic Tasks
Natural language inference and paraphrasing tasks primarily involve understanding the semantic relationships and meaning between sentences or phrases. Syntax, including word order and grammatical rules, can influence the overall coherence and clarity of the expressions, but it is not the primary focus of these tasks.
The results depicted in Figures 5 and 6 illustrate the impact of nullifying PE in tasks related to natural language inference and paraphrasing. We notice a consistent pattern emerge in the graphs where morphologically poor languages are notably affected by the absence of positional encoding, while the impact is comparatively less for morphologically rich languages. However, in contrast to syntactic tasks, the variability in impact across different languages is less pronounced for semantic tasks.
Task (Metric) Correlation
NER (F1) -0.742
POS (F1) -0.693
Dependency Parsing (UAS) -0.882
Dependency Parsing (LAS) -0.873
XNLI (Accuracy) -0.773
PAWS-X (Accuracy) -0.486
Table 1: Spearman correlation coefficient between morphological complexity of a language and relative decrease in performance across different tasks
In Table 1, we report the statistical correlation between the morphological complexity of a language and the relative decrease in performance across tasks, as determined by the Spearman Correlation Coefficient. A strong negative correlation is observed, indicating that higher morphological complexity is associated with a lower relative decrease in performance.
6 Conclusion
In this work, we demonstrate the interplay between positional encoding and morphology for morphologically diverse languages. We present the first study regarding the varying impact of positional encoding across languages with varying morphological complexity. Our results reveal that the importance of positional encoding diminishes as the morphological complexity of a language increases. Our study also motivates the need for a deeper understanding of positional encoding, augmenting them to better reflect the different languages under consideration.
Ethics Statement
Limitations
Appendix A The Relationship between Morphology and Word Order
@@
Linguistic theories : Morphological case markings serve a similar function as word order (Sapir, 1921; Blake, 2001). Several theories suggest that the presence of morphological case is necessary for free word order in a language (Comrie, 1989; Haspelmath, 1999). Either morphological case or structural position facilitates the unambiguous determination of grammatical role of the constituents of a sentence. The existence of morphological case reduces the need for fixed structural position in determination of the grammatical function of a word or phrase, allowing for variable word order.sla However, if morphological case is absent, fixed placement of words (and phrases) is necessary, exhibiting a fixed or rigid word order. In this work, we align our empirical study in accordance with the above theories that hints at the existence of a correlation between morphology and word order. Specifically, morphologically rich languages which tend to exhibit higher word-order flexibility as compared to morphologically poor languages [^1] The theory regarding morphology and word order is a linguistically complex topic. Concurrent theories in the literature propose that there is no correlation between morphological complexity and word order. Müller (2002) demonstrates phenomena like scrambling and topic shift, where the change in word order does not necessarily require a high level of morphological complexity. However, this is beyond the scope of our study..
A.1 Spectrum of Morphological Complexity
@@
Appendix B Details of Languages
We provide an overview of the languages included in our study in Table 2. Additionally, Table 4 presents the details of the treebanks used in the dependency parsing experiments.
Appendix C TTR-based Morphological Complexity
The TTR-based morphological complexity of different languages are listed in Table 3. For space-delimited languages, we use the tokenizer from NLTK library for word segmentation.
@@ Appendix C TTR-based Morphological Complexity
Basque (eu) Basque orai-nlp/ElhBERTeu
Bengali (bn) Indo-European: Indo-Aryan sagorsarker/bangla-bert-base
@@
Dutch (nl) Indo-European: Germanic GroNLP/bert-base-dutch-cased
@@
Hungarian (hu) Uralic SZTAKI-HLT/hubert-base-$c$
@@
Language (ISO code) FLORES-200 code TTR
Arabic (ar) arb_Arab 0.359
Basque (eu) eus_Latn 0.324
Bengali (bn) ben_Beng 0.292
Bulgarian (bg) bul_Cyrl 0.268
Chinese (zh) zho_Hans 0.17
Dutch (nl) nld_Latn 0.207
English (en) eng_Latn 0.194
Finnish (fi) fin_Latn 0.428
French (fr) fra_Latn 0.191
German (de) deu_Latn 0.244
Greek (el) ell_Grek 0.253
Hebrew (he) heb_Hebr 0.364
Hungarian (hu) hun_Latn 0.345
Indonesian (id) ind_Latn 0.195
Italian (it) ita_Latn 0.217
Korean (ko) kor_Hang 0.465
Portuguese (pt) por_Latn 0.205
Russian (ru) rus_Cyrl 0.334
Spanish (es) spa_Latn 0.192
Swahili (sw) swh_Latn 0.212
Turkish (tr) tur_Latn 0.376
Vietnamese (vi) vie_Latn 0.077
Table 3: TTR-based morphological complexity of different languages
Language Treebank
Chinese (zh) UD_Chinese-GSD
Portuguese (pt) UD_Portuguese-Bosque
Spanish (es) UD_Spanish-GSD
English (en) UD_English-GUM
French (fr) UD_French-GSD
Italian (it) UD_Italian-ISDT
Russian (ru) UD_Russian-Taiga
German (de) UD_German-GSD
Basque (eu) UD_Basque-BDT
Finnish (fi) UD_Finnish-FTB
Turkish (tr) UD_Turkish-Penn
Table 4: Details of treebanks of different languages
Appendix D Additional Results
The effect of removing positional encoding in dependency parsing is examined by analyzing the relative decrease in UAS (Figure 4) and LAS scores (Figure 7).
Figure 7: Effect of Positional Encoding on Dependency Parsing (LAS).
Appendix E Results on GLUE tasks